A fruit company delivers its fruit in two types of boxes: large and small. A delivery of 5 large boxes and 3 small boxes has a total weight of 95 kilograms. A delivery of 2 large boxes and 6 small boxes has a total weight of 68 kilograms. How much does each type of box weigh?
Answer by greenestamps(13200)



Imagine another delivery half of the second one shown -- 1 large box and 3 small boxes with a weight of 34kg.

Now compare that to the first delivery shown. The difference is 4 large boxes, with a difference of 61kg in weight. So the weight of each large box is 61/4 kg.

Use that in any of the three deliveries to determine the weight of each small box.

With formal algebra, the solution looks virtually the same:

5l+3s=95
2l+6s=68
l+3s=34
4l=61
l=61/4

61/4+3s = 34
...